OpenStudy (anonymous):

Find the maximum or minimum of the following quadratic function: y =x2 - 10x + 25. A. 0 B. 25 C. -25 D. 50

OpenStudy (anonymous):

differentiation? maxima and minima?

OpenStudy (campbell_st):

this is a perfect square, y = (x - 5)^2 it contacts the x axis at 1 point, which will be the minimum. all you need to do is solve x - 5 = 0 to find the point of contact... substitute to find y, the minimum value.

OpenStudy (anonymous):

\[\frac{ dy }{ dx} = 2x - 10\] when you equal the first derivative with 0 you get: 2x -10 = 0 and x = 5 your function has the minimum value in x = 5 and the value is y = 25 - 50 + 25 = 0